# Apollon Real-Time Adaptive Optics (ARTAO) - Astronomy-Inspired Wavefront Stabilization in Ultraintense Lasers

### v1

@@ 6.3 Short-term Stabilization
Figure 12: Top: time series measurement of the calculated Strehl ratio (compared to the reference WF) of the pilot beam, where ARTAO is activated at t=0. Bottom: the histogram of the corresponding data series.
@@
In these experiments, we mainly adjusted the framerate of the WFS, the Tip/Tilt gain and the modal gain curves. For each setting, we incrementally increased the feedback gain to the edge of stability. We then recorded a time sequence of WFs of one minute prior to closing the loop and continued to record for another minute. From the WFs, we calculated the Strehl ratio, which we used as relative measure for the beam stability. This Strehl ratio obviously only applies if all static aberrations are removed up to the target area. The result is shown in Fig. 12. The loop exhibits excellent performance on this short time scale, increasing the minimum Strehl ratio from 0.65 to 0.96.
@@
Figure 14: Time series measurement of the beam pointing of the pilot beam, where ARTAO is activated at t=0. The inserts show a zoomed in portion to illustrate the fast oscillations in the beam pointing.
@@ 7.1 Long-Term Stability Issues
These instability events pose a significant risk for operation, as large DM strokes during these phases can create hotspots in the beam, potentially damaging system optics during shot delivery. The unstable wavefronts exhibit wave-like structures (Fig. 16, right inset) oscillating over time, strongly suggesting lateral misregistration between the WFS and DM30. This is likely due to environmental changes deforming the beamline slightly and thus shifting the DM image on the WFS over time.
This hypothesis was confirmed when minor adjustments to a steering mirror outside the DM conjugation plane restored stability. Thus, implementing a tracking routine to maintain lateral beam alignment on the WFS is a promising countermeasure. This approach will require a sensitive analysis routine to detect beam drift early 30.
@@ 7.2 Nearfield Intensity Variations
Figure 17: Imprints of artificially large DM strokes onto the NF fluence in a non-conjugate image plane.
Another issue arises from the coupling between the pre-compensated WF and the NF intensity. Aberrations introduced in planes not conjugated to the DM lead to pre-compensation effects that couple the WF to NF intensity variations. Depending on the magnitude of the aberrations, this coupling can generate hotspots, reducing the system’s maximum supportable pulse energy. Artificial examples demonstrating this effect are shown in Fig. 17, where deliberate large DM strokes, approx. one order of magnitude higher than in regular operation, illustrate the potential for hotspot generation. Such amplitudes do occur for unstable loops.
This coupling is more pronounced for smaller beam diameters due to stronger WF gradients for the same phase amplitude. Mitigation strategies therefore should focus on reducing pre-compensation stroke magnitudes. Assuming passive aberration mitigation strategies are optimized, two potential solutions emerge:
@@
Both approaches require further investigation in future work.

### v2

@@ 6.3 Short-term Stabilization
Figure 12: Top: time series measurement of the Strehl ratio (compared to the reference WF, calculated via the FFT of the measured NF) of the pilot beam, where ARTAO is activated at t=0. Bottom: the histogram of the corresponding data series.
@@
In these experiments, we mainly adjusted the framerate of the WFS, the Tip/Tilt gain and the modal gain curves. For each setting, we incrementally increased the feedback gain to the edge of stability. We then recorded a time sequence of WFs of one minute prior to closing the loop and continued to record for another minute. From the WFs, we calculated the Strehl ratio via a FFT of the measured NF, which we used as relative measure for the beam stability. This Strehl ratio obviously only applies if all static aberrations are removed up to the target area and is thus an unrealistic value from the experiment point of view, but it remains useful to build some intuition on the performance of the AO loop in terms of achievable stability. The result is shown in Fig. 12. The loop exhibits excellent performance on this short time scale, increasing the minimum Strehl ratio from 0.65 to 0.96.
@@
Figure 14: Time series measurement of the beam pointing of the pilot beam, where ARTAO is activated at t=0. The tilt X (red) and tilt Y (blue) curves represent the tilt-portion of the recorded WF, relative to the reference, and is given in peak-to-valley in terms of the central wavelength of Apollon, roughly corresponding to the movement of the focal spot in focal spot diameters. The inserts show a zoomed in portion to illustrate the fast oscillations in the beam pointing.
@@ 7.1 Long-Term Stability Issues
These instability events pose a significant risk for operation, as large DM strokes during these phases can create hotspots in the beam, potentially damaging system optics during shot delivery. The unstable wavefronts exhibit wave-like structures (Fig. 16, right inset) oscillating rapidly over time, strongly suggesting lateral misregistration between the WFS and DM30. This is likely due to environmental changes, e.g. temperature or humidity deforming the beamline slightly and thus shifting the DM image on the WFS over time.
This hypothesis was confirmed when minor manual adjustments to a randomly picked steering mirror outside the DM conjugation plane between the DM and the WFS realigned the DM image position, which in turn restored stability. Thus, implementing a tracking routine to maintain lateral beam alignment on the WFS is a promising countermeasure. This approach will require a sensitive analysis routine to detect beam drift early 30.
It should be noted, however, that other sources of instability are still worth consideration. This includes, e.g., a drop in illumination on the WFS, software crashes and exceedingly fast vibrations. This implies that a robust machine safety framework is necessary in any case.
@@ 7.2 Nearfield Intensity Variations
Figure 17: Imprints of artificially large DM strokes onto the NF fluence in a non-conjugate image plane, where each NF corresponds to a different set of random actuator positions.
Another issue arises from the coupling between the pre-compensated WF and the NF intensity. Aberrations introduced in planes not conjugated to the DM lead to pre-compensation effects that couple the WF to NF intensity variations. Depending on the magnitude of the aberrations, this coupling can generate hotspots, reducing the system’s maximum supportable pulse energy. Artificial examples demonstrating this effect are shown in Fig. 17, where deliberate large DM strokes, approx. one order of magnitude higher than in regular operation, illustrate the potential for hotspot generation. Such amplitudes do occur for unstable loops, but could also occur for large amounts of pre-compensation, even if the loop remains stable. In this case, experiments may still be feasible at reduced energy, while the impact of the distorted NF and thus the three-dimensional intensity distribution around the focal plane on the target interaction has to be considered for each experiment individually.
The WF to NF coupling is more pronounced for smaller beam diameters due to stronger WF gradients for the same phase amplitude. Mitigation strategies therefore should focus on reducing pre-compensation stroke magnitudes. Assuming passive aberration mitigation strategies are optimized, two potential solutions emerge:
@@
Both approaches require further investigation in future work, but in general, reducing and possibly eliminating the pre-compensation amplitude as far as possible is the favorable path.